Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2015-10-03

Observers: Joe Lyman, Heather Campbell
TAT: Joe Anderson, Cosimo Inserra
Data reducers: Morgan Fraser

Pointing restrictions for most of the night, limiting observations to southern targets. Seeing also quite poor preventing some critical photometry requiring good conditions. Intermittent clouds at start and end of the night. EFOSC used for the whole night.

UT      Target                  Comment
23:30                           Dome open, pointing restriction south
23:30   LTT7379                 Spec standard
23:55   OGLE15gt                Classification (thin clouds, not useful)
00:10   LSQ15adm                Photometry BVgri
00:20                           Seeing ~1.4"
00:30   ASASSN-15oz             Follow up 900s gr13
00:55   OGLE15gt                Repeat Classification (AGN/galaxy?)
01:20   ASASSN-15oi             Follow up 2x2700s gr11 & gr16
03:40                           Seeing ~1.5"
04:55   OGLE15ic                Classification slit 1.5"
05:15   OGLE15jy                Classification (noisy, Ib/IIb?)
05:15                           Wind 14m/s
05:30                           Thin clouds beginning to appear
05:45   SN2015F                 Follow up gr16 - abandoned after first exposure - clouds
06:40   LSQ15bco                Classification (young IIP?)
07:10   OGLE15kk                Classification (Ia)
07:45                           Pointing restriction removed
07:45   PSNJ02051332+0606084    Follow up 1000s gr11 & gr16
08:35   PS15cbm                 Classification (Ia post max)
09:15   L745a                   Spec standard